a body covers half of its total height in the last second of its free fall calculate i)duration  of fall ii)height from which body was dropped


 

Grade:

5 Answers

Aman Bansal
11 years ago

Dear Student,

Time of journey (free fall) = t = n seconds
Height of free fall = H meters ; 
initial velocity = 0 ; acceleration due to gravity = g = 9.8 m/s²
Hn = (1/2) g t² = (1/2) g n² = distance travelled in n seconds
H(n-1) = (1/2) g (n - 1)² = distance travelled in n - 1 seconds
Distance travelled in the n th (last) second = h = Hn - H(n-1)
= (1/2)g {n² - (n - 1)²} = (1/2) g (2n - 1)
As per the question : h = (1/2) Hn
(1/2) g (2n - 1) = (1/4) g n²
=> 2n - 1 = n²/2
=> n² = 4n - 2
=> n² - 4n + 2 = 0
=> (n - 2)² - 2 = 0 
=> (n - 2)² = 2
=> n - 2 = ± √2
=> n = 2 ± √2
=> n = (2 + √2) or (2 - √2)
n = 2 - √2, being less than 1, is rejected.
Time of journey = n = 2 + √2 seconds
Height of fall = (1/2) g (2 + √2)² = (1/2)(9.8)(4 + 2 + 4√2) = 9.8*(3 + 2√2) meters
(Calculate by taking √2 = 1.414

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

http://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert


Afzaal
7 years ago
A body is failling from rest at a certain hight.The distance covered by body in first 3 second is equal to the distance coverd by body in last second before striking the ground.How long body remains in air
Kajal singh
6 years ago
T2=2h/g: (3)Square=2h/10: h=45mS=u+1/2a(2n-1) ((foumula to find displacement in one second))-45=0+1/2*-10(2t-1) where n= tseconds(2t-1) =45*2/10T=5seconds
mahesh kumar
4 years ago
u=0
half distance in last second
[h/2= g/2[2n-1] ….....1
fall freely from rest
h= 1/2gn^2 ….........2
from equation 1and 2 we find
g/2[2n-1]=g/4n^2
2n-1=n^2/2
4n-2=  n^2
n^2-4n+2=0
[n-2]^2-2 =0 
[n-2]^2 =2
n-2 =2^1/2
n=2^1/2+2 .
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
Time of journey (free fall) = t = n seconds
Height of free fall = H meters ; 
initial velocity = 0 ; acceleration due to gravity = g = 9.8 m/s²
Hn = (1/2) g t² = (1/2) g n² = distance travelled in n seconds
H(n-1) = (1/2) g (n - 1)² = distance travelled in n - 1 seconds
Distance travelled in the n th (last) second = h = Hn - H(n-1)
= (1/2)g {n² - (n - 1)²} = (1/2) g (2n - 1)
As per the question : h = (1/2) Hn
(1/2) g (2n - 1) = (1/4) g n²
=> 2n - 1 = n²/2
=> n² = 4n - 2
=> n² - 4n + 2 = 0
=> (n - 2)² - 2 = 0 
=> (n - 2)² = 2
=> n - 2 = ± √2
=> n = 2 ± √2
=> n = (2 + √2) or (2 - √2)
n = 2 - √2, being less than 1, is rejected.
Time of journey = n = 2 + √2 seconds
Height of fall = (1/2) g (2 + √2)² = (1/2)(9.8)(4 + 2 + 4√2) = 9.8*(3 + 2√2) = 57.12 m
 
Thanks and regards,
Kushagra